OpenStudy (imtiaz7):

Express Log-2logx+3log (x+1) -log (x^2 -1) as a single logarithm

OpenStudy (whpalmer4):

With a few properties of logarithms you should be able to combine that into a single logarithm: \[\log a+\log b = \log a*b\]\[\log a - \log b = \log \frac{a}{b}\]\[\log x^a = a\log x\]

OpenStudy (imtiaz7):

thanks whpalmer4

OpenStudy (solomonzelman):

\(\large\color{blue}{ \sf Log(-2)~log(x)+3log (x+1) -log (x^2 -1)}\)

OpenStudy (solomonzelman):

(hate disconnections)

OpenStudy (whpalmer4):

Is that supposed to be \[-2\log x + 3\log(x+1) - \log(x^2-1)\]? I'm confused by the initial "Log-"...you can't take the logarithm of a negative number if you expect a real answer.

OpenStudy (solomonzelman):

( without missing the 1st 3 letters of the questions in the blue box ) \(\large\color{blue}{ \sf Log(-2)~log(x)+3log (x+1) -log (x^2 -1)}\)

OpenStudy (solomonzelman):

it can't be just log, so I guess the question is what I wrote, starting from log(-2)

OpenStudy (whpalmer4):

Tell me, what power do you raise 10 to to get a negative number? :-)

OpenStudy (whpalmer4):

\[\log_b x =a\]means that \[b^a=x\]\[10^a=-2\]\[a=\]

OpenStudy (solomonzelman):

https://www.google.com/#q=log(-2) WOW holy cow, that's weird!

OpenStudy (solomonzelman):

idk

OpenStudy (solomonzelman):

@Imtiaz7 can you re-write the question as clearly as possible please ?

OpenStudy (whpalmer4):

The natural log is more, well, natural: \[\ln -2 = i\pi + \ln 2 \approx 0.693147+3.14159i\] Your manipulation wasn't quite correct. \[\log_{10}x = -2\rightarrow 10^{-2} = x = \frac{1}{10^2} = 0.01\] But that's not the problem I posed to you. \[\log_{10} -2 = x\rightarrow 10^x = -2\]

OpenStudy (imtiaz7):